Given that f(x) = 2x^2 - x + 5 , find
a) f(-1)
b) f(a+2)
Thanks so much in advance:)

Answer by jim_thompson5910(35256)

Given that f(x) = 2x^2 - x + 5 , find
a) f(-1)
b) f(a+2)
Thanks so much in advance:)

----------------------------------------------------------------------

a)

Plug in x = -1 and evaluate.

f(x) = 2x^2 - x + 5

f(-1) = 2(-1)^2 - (-1) + 5 ... replace ALL copies of x with -1 (since x = -1)

f(-1) = 2(1) - (-1) + 5

f(-1) = 2 + 1 + 5

f(-1) = 8

-----------------

b)

Plug in x = a+2 and simplify.

f(x) = 2x^2 - x + 5

f(a+2) = 2(a+2)^2 - (a+2) + 5 ... replace ALL copies of x with a+2 (since x = a+2)

f(a+2) = 2(a+2)(a+2) - (a+2) + 5

f(a+2) = 2(a^2+4a+4) - (a+2) + 5

f(a+2) = 2a^2+8a+8 -a-2 + 5

f(a+2) = 2a^2+7a+11